### v2

# DP-CDA: An Algorithm for Enhanced Privacy Preservation in Dataset Synthesis Through Randomized Mixing

Abstract
In recent years, the growth of data across various sectors, including healthcare, security, finance, and education, has created significant opportunities for analysis and informed decision-making. However, these datasets often contain sensitive and personal information, which raises serious privacy concerns. Protecting individual privacy is crucial, yet many existing machine learning and data publishing algorithms struggle with high-dimensional data, facing challenges related to computational efficiency and privacy preservation. To address these challenges, we introduce an effective data publishing algorithm DP-CDA. Our proposed algorithm generates synthetic datasets by randomly mixing data in a class-specific manner, and inducing carefully-tuned randomness to ensure formal privacy guarantees. Our comprehensive privacy accounting shows that DP-CDA provides a stronger privacy guarantee compared to existing methods, allowing for better utility while maintaining strict level of privacy. To evaluate the effectiveness of DP-CDA, we examine the accuracy of predictive models trained on the synthetic data, which serves as a measure of dataset utility. Importantly, we identify an optimal order of mixing that balances privacy guarantee with predictive accuracy. Our results indicate that synthetic datasets produced using the DP-CDA can achieve superior utility compared to those generated by traditional data publishing algorithms, even when subject to the same privacy requirements.
keywords:
Differential privacy , Deep Learning , Dataset synthesis , Private datasets
journal: Elsevier Array
[…]
[inst1]organization=Department of Electrical and Electronic Engineering,addressline=Bangladesh University Engineering Technology, city=Dhaka, postcode=1205, country=Bangladesh
[…]
[inst5]organization=School of Data and Sciences,addressline=BRAC University, city=Dhaka, postcode=1212, country=Bangladesh
1 Introduction
In the realm of machine learning or deep learning applications, it is widely acknowledged that the dataset is of utmost importance, as it contains sensitive information like medical records, personal photos/information, or proprietary data during the training process. Organizations, including private companies, government agencies, and hospitals, routinely handle substantial volumes of personal information about their clients, customers, or patients. This data is often private or sensitive in nature, underscoring the importance of implementing systems and methodologies that can effectively analyze the data while upholding privacy and ensuring the confidentiality of individuals. It should be noted that using standard machine learning techniques to train a model can potentially lead to data breaches [1]. Therefore, a prudent approach involves considering the creation of a “distilled version” of the original dataset, and subsequently training the model exclusively using that version to ensure the protection of sensitive information [2, 3, 4]. It should be noted here that only anonymization can not provide any privacy in the presence of auxiliary information [5]. More specifically, it’s important to be aware of “composition attacks”, where malicious actors can combine algorithm output with other information to identify individuals in the dataset [6].
As a potential solution, a cryptographically motivated approach – differential privacy, has gained traction in privacy challenges. It quantifies the risk of privacy breaches using a parameter $\epsilon$, which governs the extent to which the outcome of a private algorithm can vary when an individual’s data is included or excluded from the dataset. A smaller $\epsilon$ value reduces the ability of adversaries to infer information about individuals in the dataset, enhancing overall privacy protection. The concept of differential privacy, introduced in [7] has gained widespread acceptance in various research fields. While numerous methods for differentially private data publishing have been proposed, many are not suitable for high-dimensional data due to their significant computational demands [8] and utility penalties. This emphasizes the importance of developing more efficient algorithms capable of handling large, complex datasets while still ensuring privacy protection.
Figure 1: Differentially private datasets generated from the MNIST, FashionMNIST, and CIFAR10 datasets, each incorporating varying order of mixture ($l$). The parameter $\epsilon$ refers to the differential privacy level; a lower value of $\epsilon$ signifies a greater degree of privacy protection.
Figure 2: Utility as a function of the order of mixture $l$.
Motivated by this, we present a synthetic privacy-preserving data publishing algorithm – Differentially Private Class-Centric Data Aggregation (DP-CDA), that integrates several data samples from the same class. More specifically, we (uniformly) randomly select $l$ data samples from a specific class, and then induce Gaussian noise (parameterized by variance terms $\sigma_{x}$, $\sigma_{y}$) to generate each synthetic data sample. This distilled dataset is subsequently utilized in the conventional training process within deep learning frameworks. Few samples of the synthetic data samples for different algorithm parameters are shown in Figure 1, whereas detailed results are presented in Section 4. We summarize our main contributions below:
  • 1.
    We analyze the synthetic data generation algorithm to achieve a tighter overall privacy guarantee.
  • 2.
    We investigate the effect of order of mixture $l$ on the performance of the model trained on the synthetic data. We empirically demonstrate that for a given dataset and privacy level, there exists an optimal $l^{*}$ for which the model performance peaks, as illustrated in the toy plot in Figure 2.
  • 3.
    Lastly, through our tighter theoretical analysis and experimental validation, we demonstrate that the synthetic data generation algorithm can achieve a given level of privacy while ensuring better accuracy compared to existing approaches.
Related Works. Over the years, numerous algorithms have been thoughtfully developed for differentially private data publishing [9, 10]. Zhu et al. conducted a comprehensive analysis of these algorithms, offering a clear and detailed overview [3]. However, these algorithms face challenges when scaling with larger data sets. That is, the computational complexity increases significantly with high-dimensional data, often exhibiting exponential growth [8]. As a result, many of these algorithms can become impractical for deep learning applications, particularly those that involve high-dimensional datasets.
Local Perturbation is a widely adopted algorithm for data publishing within the framework of differential privacy [11, 12, 13, 14]. This technique involves the addition of noise from a specific distribution to data points, allowing the generated synthetic data to remain consistent with the original data domain. This compatibility makes it suitable for use with existing learning algorithms. However, to ensure robust privacy, noise with high variance must be introduced, which adversely affects model performance. In contrast, random projection reduces high-dimensional data to a lower-dimensional subspace prior to inducing noise [15, 16]. This method results in a new data domain, which can limit the applicability of models that are specific to a given domain.
There is a considerable amount of work on the application of differentially private algorithms to various machine learning problems, including regression [17, 18, 19], online learning [20], graphical models [21], empirical risk minimization [22, 23, 24, 22, 5, 25], and deep learning [26, 27, 28]. The objectives of these algorithms differ slightly from those of the data publishing algorithms mentioned earlier. While differentially private data publishing algorithms focus on designing differentially private datasets at the input end of the pipeline, differentially private algorithms aim to create learning algorithms that yield differentially private models at the output end. Any model trained on a dataset that is differentially private will also maintain differential privacy. [7], the goals of these learning algorithms represent a more relaxed version of the data publishing algorithms. Unfortunately, these differentially private algorithms do not always guarantee data privacy in adverse scenarios. If an attacker gains access to the input side of the pipeline, the entire algorithm may be compromised, thus undermining the security of user data.
In the past few years, several methods have been proposed for differentially private data publishing algorithms using mixtures. Karakus et al. illustrated that it is possible to effectively train a straightforward linear model by utilizing mixtures that are free from noise [29]. The exploration of learning nonlinear models with mixtures has also been addressed in previous studies [30, 31, 32, 33, 34]. For instance, Tokozune et al. demonstrated that sound recognition models can be trained using mixtures of audio signals [30]. Similarly, other studies have indicated that it is feasible to train an image classification model using image mixtures [31, 32, 33]. However, these studies primarily focused on noiseless mixtures involving only two or three data points. Note that, Lee et al. eventually proved that it is possible to train deep neural networks with a high order of mixture while incorporating added noise to enhance privacy [8].
Recently, data publishing algorithms have been automated by incorporating a deep learning framework  [35]. The integration of a generative adversarial network for producing differentially private datasets has gained much popularity  [36],  [37]. Although GAN-generated synthetic images offer differential privacy (DP) guarantees and maintain the utility of the original private images, they often appear visually similar to those private images. This similarity poses a challenge in fulfilling a key privacy requirement, which is the need for visual dissimilarity. Furthermore, training a GAN can be difficult due to several issues, including vanishing gradients, mode collapse, training instability, and convergence failures. In this work, we consider the problem of releasing synthetic data for training an image classification model, while ensuring $(\epsilon,\delta)$-Differential Privacy (DP) guarantees and providing utility close to that of conventional DP training. We note that the synthetic data-generating mechanism is computationally simple, and can be adapted for high-dimensional data.
2 Background and Problem Formulation
We reviewed some definitions, theorems, and propositions, which are necessary for our problem formulation, according to [5, 38, 25] in the supplementary materials (A).
Notation. We use lower-case bold-faced letters (e.g., $\mathbf{x}$) for vectors, upper-case bold-faced letters (e.g., $\mathbf{X}$) for matrices, and unbolded letters (e.g., $N$ or $n$) for scalars. The set of $n$ data points $(\mathbf{x}_{i},y_{i})$ is represented as $\mathbb{D}=\{(\mathbf{x}_{1},y_{1}),(\mathbf{x}_{2},y_{2}),\ldots,(\mathbf{x}_{n},y_{n})\}\triangleq\{x_{i},y_{i}\}_{i=1}^{n}$. We use $\|\cdot\|_{2}$ and $\|\cdot\|_{F}$ to denote the $\ell_{2}$-norm of a vector and the Frobenius norm of a matrix.
Problem Formulation. As mentioned before, numerous modern ML algorithms are trained on privacy-sensitive data, and the model parameters can cause significant privacy leakage. In this work, we focus on generating synthetic data that satisfies formal and strict privacy guarantees that can be utilized to train the aforementioned ML models. More specifically, given a dataset $\mathbb{D}=\{(x_{i},y_{i})\}_{i=1}^{n}$, where each $\mathbf{x}_{i}\in\mathbb{R}^{d_{x}}$ is a feature vector and $y_{i}$ is a label, the goal is to construct a new dataset $\mathbb{D}^{\prime}$ that ensures privacy, while maintaining high utility for downstream ML tasks. The new dataset $\mathbb{D}^{\prime}$ should enable effective training of a machine-learning model $f(\mathbf{w})$, yielding predictive performance comparable to that achieved with the original dataset $\mathbb{D}$. This requirement can be formulated as a utility constraint:
$\text{Utility}(f,\mathbb{D}^{\prime})\geq\theta\text{Utility}(f,\mathbb{D}),$
where, $\text{Utility}(f,\mathbb{D}^{\prime})$ is a metric (e.g., accuracy) that reflects the performance of model $f(\mathbf{w})$ when trained on $\mathbb{D}^{\prime}$, and $\theta\in(0,1]$ is a predefined utility threshold indicating acceptable performance relative to training on $\mathbb{D}$. As mentioned before, our objective of this work is to design a transformation mechanism $\mathcal{M}$ that maximizes privacy protection while simultaneously meeting the utility constraint for effective model performance. This can be expressed as:
$\text{minimize}\quad\epsilon$
subject $\text{subject to}\quad\text{Utility}(f,\mathbb{D}^{\prime})\geq\theta\text{Utility}(f,\mathbb{D}).$
3 Proposed Privacy-Preserving Synthetic Data Generation Algorithm (DP-CDA)
In this paper, we describe the DP-CDA – a privacy-preserving synthetic data generation algorithm according to [8]. Let us assume we have a dataset $\mathbb{D}=\{(x_{i},y_{i})\}_{i=1}^{N}$. We can form the design matrix $\hat{\mathbf{X}}\in\mathbb{R}^{N\times d_{x}}$ and $\mathbf{y}\in\mathbb{R}^{N}$, where the $i$-th sample $\mathbf{x}_{i}\in\mathbb{R}^{d_{x}}$ is the $i$-th row in $\hat{\mathbf{X}}$, and $y_{i}\in\{1,2,3,...,K\}$ is the corresponding class label. We intend to synthesize $T$ new data points by mixing $l$ data points from the same class. We first perform a feature-wise $z$-score normalization of the data points $\mathbf{x}_{i}$ as
$x_{ij}=\frac{x_{ij}-\mu_{j}}{\sigma_{j}}\quad\forall j\in\{1,2,\ldots,d_{x}\},$
where $x_{ij}$ is the $j$-th entry of the $d_{x}$-dimensional feature vector $\mathbf{x}_{i}$, $\mu_{j}=\frac{1}{N}\sum_{i=1}^{N}x_{ij}$, and $\sigma_{j}=\sqrt{\frac{1}{N}\sum_{i=1}^{N}(x_{ij}-\mu_{j})^{2}}$. Next, we normalize each sample $\mathbf{x}_{i}$ with respect to its $\ell_{2}$-norm as follows:
$\mathbf{x}_{i}=\frac{\mathbf{x}_{i}}{\max(1,\|\mathbf{x}_{i}\|_{2}/c)},$
where $c$ is the $\ell_{2}$ norm clipping parameter [27]. We denote the design matrix with the preprocessed data samples with $\mathbf{X}$. Note that, the labels $y_{i}$ are converted to one-hot encoded vectors, and the labels matrix is denoted as $\mathbf{Y}\in\mathbb{R}^{N\times K}$. Our goal now is to generate a synthetic dataset $\tilde{\mathbf{X}}\in\mathbb{R}^{T\times d_{x}}$ and corresponding one-hot-encoded labels $\tilde{\mathbf{Y}}\in\mathbb{R}^{T\times K}$. Since there are $K$ classes, we intend to ensure that each class of the synthetic dataset will have approximately $T_{k}=\left\lfloor\frac{T}{K}\right\rfloor$ synthetic samples. For each class $k\in\{1,2,\dots,K\}$, we uniformly randomly select $l$ data points (order of mixture) without replacement from the subset $\tilde{\mathbf{X}}_{k}=\left\{\mathbf{x}_{i}:y_{i}=k\right\}$. Let $\left\{\mathbf{x}_{i_{1}},\mathbf{x}_{i_{2}},...,\mathbf{x}_{i_{l}}\right\}$ be the randomly selected data points from $\mathbf{X}_{k}$. A synthetic sample $\tilde{\mathbf{x}}_{t}^{(k)}$ is generated as:
$\tilde{\mathbf{x}}_{t}^{(k)}=\frac{1}{l}\sum_{j=1}^{l}\mathbf{x}_{i_{j}}+\mathbf{n}_{x},\ \forall t\in\{1,2,\ldots,T_{k}\},$
where $\mathbf{n}_{x}\sim\mathcal{N}(0,\sigma_{x}^{2}\mathbf{I}_{d_{x}})$ is a $d_{x}$-dimensional Gaussian random variable with covariance matrix $\sigma_{x}^{2}\mathbf{I}_{d_{x}}$, and $\mathbf{I}$ is a $d_{x}\times d_{x}$ identity matrix.
Similarly, the corresponding synthetic label $\tilde{y}_{t}^{(k)}$ is computed in two steps. First, the mean of the corresponding one-hot encoded labels of the $l$ randomly selected data points are computed, and noise drawn from a Gaussian distribution is added to it as follows:
$\tilde{y}_{t}^{\text{one-hot(k)}}=\frac{1}{l}\sum_{j=1}^{l}y_{i_{j}}^{\text{one-hot}}+\mathbf{n}_{y},$
where $\mathbf{n}_{y}\sim\mathcal{N}(0,\sigma_{y}^{2}\mathbf{I}_{K})$ is a $K$-dimensional Gaussian random variable with covariance matrix $\sigma_{y}^{2}\mathbf{I}_{K}$, $\mathbf{I}$ is a $K\times K$ identity matrix, and $y_{i_{j}}^{\text{one-hot}}$ is the one-hot-encoded label vector corresponding to $\mathbf{x}_{i_{j}}$. Then these one-hot encoded labels are then converted to integer labels as follows:
$\tilde{y}_{\text{t}}^{\text{(k)}}=\arg\max_{i\in\{0,1,\dots,K-1\}}\tilde{y}_{\text{t}}^{\text{one-hot(k)}}[i].$ (1)
This process is repeated $T_{k}$ times for each class $k$ to form a class-balanced synthetic dataset $(\tilde{\mathbf{X}},\tilde{\mathbf{y}})$. The complete synthetic dataset generation algorithm is shown in Algorithm 1. It is easy to show that the computational complexity of the proposed algorithm is $\mathcal{O}(Tld_{x})$.
Algorithm 1 DP-CDA: Privacy-Preserving Synthetic Data Generation
0:  Dataset $\mathbb{D}=\{(\mathbf{x}_{i},y_{i})\}_{i=1}^{N}$, number of synthetic samples $T$, order of mixture $l$, $\ell_{2}$ norm clipping parameter $c$, noise parameters $\sigma_{x}$ and $\sigma_{y}$
0:  Synthetic dataset $(\tilde{\mathbf{X}},\tilde{\mathbf{Y}})$
1:  Formulate the design matrix $\hat{\mathbf{X}}\in\mathbb{R}^{N\times d_{x}}$ and label vector $\mathbf{y}\in\mathbb{R}^{N}$
2:  for each feature $j\in\{1,2,\ldots,d_{x}\}$ do
3:     Compute $\mu_{j}=\frac{1}{N}\sum_{i=1}^{N}x_{ij}$ and $\sigma_{j}=\sqrt{\frac{1}{N}\sum_{i=1}^{N}(x_{ij}-\mu_{j})^{2}}$
4:     Compute $x_{ij}=\frac{x_{ij}-\mu_{j}}{\sigma_{j}}$ $\forall i\in\{1,2,\ldots,N\}$
5:  end for
6:  Normalize $\mathbf{x}_{i}=\frac{\mathbf{x}_{i}}{\max(1,\|\mathbf{x}_{i}\|_{2}/c)}$ $\forall i\in\{1,2,\ldots,N\}$
7:  Convert $\mathbf{y}$ to one-hot encoded vectors $\mathbf{Y}\in\mathbb{R}^{N\times K}$, and set $T_{k}=\left\lfloor\frac{T}{K}\right\rfloor$
8:  for each class $k\in\{1,2,\ldots,K\}$ do
9:     for each synthetic sample $t\in\{1,2,\ldots,T_{k}\}$ do
10:        Uniformly randomly select $l$ data points $\{\mathbf{x}_{i_{1}},\mathbf{x}_{i_{2}},\ldots,\mathbf{x}_{i_{l}}\}$ from $\{\mathbf{x}_{i}:y_{i}=k\}$
11:        Compute $\tilde{\mathbf{x}}_{t}^{(k)}=\frac{1}{l}\sum_{j=1}^{l}\mathbf{x}_{i_{j}}+\mathbf{n}_{x}$, where $\mathbf{n}_{x}\sim\mathcal{N}(0,\sigma_{x}^{2}\mathbf{I}_{d_{x}})$
12:        Compute $\tilde{y}_{t}^{\text{one-hot(k)}}=\frac{1}{l}\sum_{j=1}^{l}y_{i_{j}}^{\text{one-hot}}+\mathbf{n}_{y}$, where $\mathbf{n}_{y}\sim\mathcal{N}(0,\sigma_{y}^{2}\mathbf{I}_{K})$
13:        Convert to integer label $\tilde{y}_{t}^{(k)}=\arg\max_{i\in\{0,1,\ldots,K-1\}}\tilde{y}_{t}^{\text{one-hot(k)}}[i]$
14:        Store $\tilde{\mathbf{x}}_{t}^{(k)}$ in $\tilde{\mathbf{X}}$ and $\tilde{y}_{t}^{(k)}$ in $\tilde{\mathbf{Y}}$
15:     end for
16:  end for
17:  return  $(\tilde{\mathbf{X}},\tilde{\mathbf{Y}})$
Figure 3: Privacy guarantee as a function of noise parameters.
Theorem 1 (Privacy of DP-CDA (Algorithm 1)).
Consider Algorithm 1 with input (privacy-sensitive) dataset $\mathbb{D}=\{(\mathbf{x}_{i},y_{i})\}_{i=1}^{N}$, order of mixture $l$, $\ell_{2}$ norm clipping parameter $c$, and noise parameters $(\sigma_{x},\sigma_{y})$. Then Algorithm 1 releases a $(\epsilon,\delta)$-differentially private synthetic dataset $(\tilde{\mathbf{X}},\tilde{\mathbf{Y}})$ consisting $T$ samples for any $10<\delta<$ 1 and $\alpha\geq 3$, where
$\epsilon$ $=\min_{\alpha\in\{3,4,\ldots\}}T\varepsilon^{\prime}(\alpha)+\frac{\log\frac{1}{\delta}}{\alpha-1}.$
Here, $\varepsilon^{\prime}(\alpha)=\frac{1}{\alpha-1}\log\left(1+p^{2}{\alpha\choose 2}\min\left\{4(e^{\varepsilon(2)}-1),2e^{\varepsilon(2)}\right\}+4G(\alpha)\right)$, where $\varepsilon(\alpha)=\frac{\alpha}{l^{2}}\left(\frac{2c^{2}}{\sigma_{x}^{2}}+\frac{1}{\sigma_{y}^{2}}\right)$, $G(\alpha)=\sum_{j=3}^{\alpha}p^{j}{\alpha\choose j}\sqrt{B(2\left\lfloor\frac{j}{2}\right\rfloor)\cdot B(\left\lceil\frac{j}{2}\right\rceil)}$, and $B(l)=\sum_{i=0}^{l}(-1)^{i}{l\choose i}e^{(i-1)\varepsilon(i)}$.
Proof.
Consider a mechanism $\mathcal{M}$ that takes $l$ data points and is ($\alpha,\varepsilon(\alpha)$)-RDP. According to [39], a new mechanism $\mathcal{M^{\prime}}$ that selects $l$ data points out of $N$ data points uniformly at random, and applies $\mathcal{M}$ is ($\alpha,\varepsilon^{\prime}(\alpha)$)-RDP for any integer $\alpha\geq 2$, where
$\varepsilon^{\prime}(\alpha)=$
$\frac{1}{\alpha-1}\log\left(1+p^{2}{\alpha\choose 2}\min\left\{4(e^{\varepsilon(2)}-1),e^{\varepsilon(2)}\min\{2,(e^{\varepsilon(\infty)}-1)\}\right\}+4G(\alpha)\right),$
$G(\alpha)=\sum_{j=3}^{\alpha}p^{j}{\alpha\choose j}\sqrt{B(2\left\lfloor\frac{j}{2}\right\rfloor)\cdot B(\left\lceil\frac{j}{2}\right\rceil)}$, and $B(l)=\sum_{i=0}^{l}(-1)^{i}{l\choose i}e^{(i-1)\varepsilon(i)}$.
Now, computing the average of $l$ data points (Step 11 of Algorithm 1) is $(\alpha,\frac{\alpha}{2(\frac{\sigma_{x}}{\Delta_{x}})^{2}})$-RDP (Proposition 3). Similarly, computing the average of $l$ one-hot-encoded vectors (Step 12 of Algorithm 1) is $(\alpha,\frac{\alpha}{2(\frac{\sigma_{y}}{\Delta_{y}})^{2}})$-RDP (Proposition 3). Therefore, according to Proposition 2, computing a synthetic data point is $(\alpha,\varepsilon(\alpha))$- RDP, where
$\varepsilon(\alpha)$ $=\frac{\alpha}{2}(\frac{\Delta^{2}_{x}}{\sigma_{x^{2}}}+\frac{\Delta^{2}_{y}}{\sigma_{y^{2}}}).$
Since in our setup $\|\mathbf{x}_{i}\|_{2}\leq c$ and the rows of $\mathbf{Y}$ are one-hot-encoded, it is easy to show that $\Delta_{x}=\frac{2c}{l}$ and $\Delta_{y}=\frac{\sqrt{2}}{l}$. Therefore,
$\varepsilon(\alpha)$ $=\frac{\alpha}{l^{2}}(\frac{2c^{2}}{\sigma_{x}^{2}}+\frac{1}{\sigma_{y}^{2}}).$
Now, since we select $l$ data points randomly out of $N$ total data points at each step $t$ (see Step 10 of Algorithm 1), generation of the $t$-th synthetic sample is ($\alpha,\varepsilon^{\prime}(\alpha)$)-RDP [39], where $\varepsilon^{\prime}(\alpha)$ is given above. Simple algebra leads to the expression $\varepsilon^{\prime}(\alpha)=\frac{1}{\alpha-1}\log\left(1+p^{2}{\alpha\choose 2}\min\left\{4(e^{\varepsilon(2)}-1),2e^{\varepsilon(2)}\right\}+4G(\alpha)\right)$. Since there are $T$ total synthetic samples are generated, the overall algorithm is $(\epsilon,\delta)$-differentially private for any $10<\delta<$ 1 according to Proposition 1, where $\epsilon=\min_{\alpha\in\{3,4,\ldots\}}T\varepsilon^{\prime}(\alpha)+\frac{\log\frac{1}{\delta}}{\alpha-1}$. ∎
Remark 1.
As mentioned before, our algorithm closely follows the algorithm proposed in [8]. However, our privacy analysis is different from that presented in [8]. More specifically, the $\varepsilon(\alpha)$ as computed in [8] is $\varepsilon(\alpha)=\frac{\alpha}{2l^{2}}\left(\frac{d_{x}}{\sigma_{x}^{2}}+\frac{d_{y}}{\sigma_{y}^{2}}\right)$. Our analysis of Algorithm 1 is much tighter and avoids the ambient data dimension. As such, our privacy guarantee is stricter than the work of [8], i.e., we achieve smaller $\epsilon$ for a given utility.
Remark 2.
In Figure 3, we illustrate how the overall privacy budget $\epsilon$ varies with the noise parameters ($\sigma_{x},\sigma_{y}$) for different values of the order of mixture $l$ for a given $N=$ 60,000. This plot clearly indicates that, for a fixed noise parameter, the privacy risk increases as $l$ increases. That is, by selecting a larger value of $l$, a smaller amount of noise will suffice to achieve a given level of privacy guarantee. Additionally, as depicted in Figure 2, the utility depends on $l$ as well. As such, there exists an optimal value of $l$ for a given dataset and noise parameters that will result in the best utility and privacy guarantee. We will demonstrate this in Section 4.
4 Experimental Results
In this section, we empirically validate the performance of our algorithm and the claim that our tighter privacy analysis of Algorithm 1 provides a stricter privacy guarantee than [8] for a given utility. For generating synthetic dataset according to Algorithm 1, we consider three diverse and well-known datasets – the MNIST dataset [40], the FashionMNIST dataset [41], and the CIFAR-10 dataset [42]. As mentioned before, we intend to train a neural network with the synthetic data generated by Algorithm 1, and then evaluate the performance of the model on the test partition of the corresponding real dataset for different order of mixture $l\in\{1,2,4,\ldots,512\}$ and noise levels $(\sigma_{x},\sigma_{y})$. From Theorem 1, we note that the overall privacy budget $\epsilon$ depends on both $l$ and $(\sigma_{x},\sigma_{y})$. For the neural network model, we consider a shallow convolutional neural network (CNN). The details of the architecture of the CNN are provided in Table 1.
Table 1: Details/parameters of the model under consideration
Layer Type Parameters Output Shape
Input - (1, 28, 28)
Conv2D Filters=32, Kernel=5x5, Stride=1, Padding=2 (32, 28, 28)
ReLU Activation - (32, 28, 28)
BatchNorm2D Num Features=32 (32, 28, 28)
MaxPooling2D Pool Size=2x2, Stride=2 (32, 14, 14)
Conv2D Filters=64, Kernel=3x3, Stride=1, Padding=1 (64, 14, 14)
ReLU Activation - (64, 14, 14)
BatchNorm2D Num Features=64 (64, 14, 14)
MaxPooling2D Pool Size=2x2, Stride=2 (64, 7, 7)
Flatten - (3136)
Fully Connected (FC1) Input=3136, Output=100 (100)
ReLU Activation - (100)
Dropout Probability=0.5 (100)
Fully Connected (FC2) Input=100, Output=100 (100)
ReLU Activation - (100)
Dropout Probability=0.5 (100)
Fully Connected (Output) Input=100, Output=10 (10)
Note that, our model includes batch normalization and dropout to enhance training stability and reduce overfitting, i.e., robust for image classification tasks. We chose to implement the Adam optimizer, setting an initial learning rate of 0.001 and employed a learning rate scheduler to adjust the learning rate as needed. As mentioned before, we recorded the accuracies on the original test partition of the base dataset, which is unmixed and noise-free.
In Table 2, we present the average accuracy and standard deviation (over 10 independent runs) for each of the aforementioned base datasets with different $l$ and privacy budgets ($\epsilon$). Note that, for each $l$, we set the noise parameters $\sigma_{x}$ and $\sigma_{y}$ in such a way that we achieve the targeted $\epsilon$. The third column $\epsilon=\infty$ is the non-privacy-preserving case $(\sigma_{x}=0,\sigma_{y}=0)$. As such, $l=1$ and $\epsilon=\infty$ indicate conventional non-privacy-preserving training with the base dataset’s training partition. From this table, we can observe that as $l$ increases for the non-private case, the utility decreases. This trend is observed across all of the three base datasets that we considered.
As shown in Figure 2, there is a specific order of mixture $l^{*}$ for which the accuracy of the model reaches a maximum. In our experiments, for all the three base datasets and different $\epsilon$ levels, we found that the model obtained the best test accuracy for $l^{*}=4$.
Finally, we compare our proposed privacy analysis and classification accuracy with those of the existing approaches in Table 3. It is evident from this table that our tighter privacy analysis provides a stricter overall privacy guarantee for a given utility level.
Table 2: Average Accuracy $\pm$ Standard Deviation for Each Dataset with Different $l$ and Privacy Budgets ($\epsilon$).
Dataset $l$ $\epsilon=\infty$ $\epsilon=20$ $\epsilon=10$
MNIST 1 0.9885 ± 0.0006 0.1080 ± 0.0055 0.1081 ± 0.0055
2 0.9746 ± 0.0011 0.6722 ± 0.1153 0.7076 ± 0.0480
4 0.9462 ± 0.0031 0.7807 ± 0.0149 0.7808 ± 0.0188
8 0.9045 ± 0.0064 0.7802 ± 0.0011 0.7673 ± 0.0156
16 0.8607 ± 0.0052 0.7691 ± 0.0183 0.7585 ± 0.0165
32 0.8237 ± 0.0106 0.7577 ± 0.0156 0.6738 ± 0.1929
64 0.7962 ± 0.0091 0.7175 ± 0.0254 0.6319 ± 0.1777
128 0.7829 ± 0.0176 0.6656 ± 0.0472 0.5431 ± 0.2180
256 0.7693 ± 0.0189 0.3620 ± 0.2573 0.3403 ± 0.2010
512 0.7603 ± 0.0181 0.2597 ± 0.1351 0.1575 ± 0.0877
FashionMNIST 1 0.9064 $\pm$ 0.0017 0.3886 $\pm$ 0.1910 0.3619 $\pm$ 0.1451
2 0.8851 $\pm$ 0.0026 0.6712 $\pm$0.0122 0.6655 $\pm$ 0.0105
4 0.8454 $\pm$ 0.0033 0.6791 $\pm$ 0.0061 0.6748 $\pm$ 0.0052
8 0.7881 $\pm$ 0.0031 0.6719 $\pm$ 0.0059 0.6703 $\pm$ 0.0095
16 0.7370 $\pm$ 0.0091 0.6652 $\pm$ 0.0058 0.6653 $\pm$ 0.0093
32 0.6970 $\pm$ 0.0118 0.6542 $\pm$ 0.0159 0.6629 $\pm$ 0.0049
64 0.6579 $\pm$ 0.0162 0.6538 $\pm$ 0.0116 0.6498 $\pm$ 0.0071
128 0.6435 $\pm$ 0.0209 0.6532 $\pm$ 0.0107 0.6399 $\pm$ 0.0184
256 0.6273 $\pm$ 0.0209 0.5243 $\pm$ 0.2124 0.6203 $\pm$ 0.0146
512 0.6181 $\pm$ 0.0221 0.5966 $\pm$ 0.0149 0.4574 $\pm$ 0.1811
CIFAR10 1 0.7210 $\pm$ 0.0035 0.1001 $\pm$ 0.0000 0.1001 $\pm$ 0.0000
2 0.6816 $\pm$ 0.0057 0.1994 $\pm$ 0.0602 0.1922 $\pm$ 0.0523
4 0.6241 $\pm$ 0.0068 0.2315 $\pm$ 0.0449 0.2313 $\pm$ 0.0455
8 0.5517 $\pm$ 0.0070 0.2258 $\pm$ 0.0435 0.1828$\pm$ 0.0681
16 0.4503 $\pm$ 0.0105 0.1761 $\pm$ 0.0657 0.1860 $\pm$ 0.0580
32 0.3758 $\pm$ 0.0076 0.1938 $\pm$ 0.0542 0.2026 $\pm$ 0.0523
64 0.3376 $\pm$ 0.0108 0.2001 $\pm$ 0.0515 0.1854 $\pm$ 0.0570
128 0.3031 $\pm$ 0.0077 0.1579 $\pm$ 0.0498 0.1341 $\pm$ 0.0457
256 0.2758 $\pm$ 0.0100 0.1449 $\pm$ 0.0455 0.1390 $\pm$ 0.0378
512 0.2574 $\pm$ 0.0068 0.1146 $\pm$ 0.0299 0.1061 $\pm$ 0.0179
Table 3: Performance comparison of the proposed algorithm with existing approaches in terms of privacy and utility.
Algorithm MNIST FashionMNIST CIFAR10
$\epsilon$ = 20 $\epsilon$ = 10 $\epsilon$ = 20 $\epsilon$ = 10 $\epsilon$ = 20 $\epsilon$ = 10
Random Projection [8] 0.104 0.100 - - 0.101 -
Local Perturbation [8] 0.098 0.098 - - 0.100 -
DP-Mix [8] 0.800 0.782 -a -a 0.269 0.244
DP-MERF [43] - 0.674 - 0.617 - -
DP-CDA (Ours) 0.796 0.795 0.685 0.680 0.276 0.276
  • a
    The way the authors in [8] computed the overall $\epsilon$ resulted in numerical instability issues for this dataset. Consequently, the experiment was not reproduced for FashionMNIST, as the noise variances could not be calculated.
5 Conclusion
In this work, we propose DP-CDA, a synthetic data generation algorithm based on the works of [8], and perform a tighter accounting of the privacy guarantee. DP-CDA combines randomly selected data points from a specific class and induces additive noise for preserving privacy. We empirically demonstrate that given a particular dataset, the number of data points utilized for generating each synthetic data point has an optimal value, which offers the best utility and dictates the overall privacy budget. Through our extensive experiments, we show that DP-CDA significantly outperforms the existing data publishing algorithms, while being less computationally complex, providing a strong privacy guarantee. An interesting direction for future work would be to incorporate DP-CDA in federated settings and to perform a theoretical analysis for choosing the optimal order of mixture.
Acknowledgement
The authors express their gratitude to the authorities of the Department of Electrical and Electronic Engineering at Bangladesh University of Engineering and Technology (BUET) for their continuous support during this research.
Appendix A Relevant Definitions and Theorems
Definition 1 (($\epsilon,\delta$)-DP [44]).
An algorithm $f:\mathcal{D}\mapsto\mathcal{T}$ provides ($\epsilon,\delta$)-differential privacy (($\epsilon,\delta$)-DP) if $\Pr(f(\mathbb{D})\in\mathcal{S})\leq\delta+e^{\epsilon}\Pr(f(\mathbb{D}^{\prime})\in\mathcal{S})$ for all measurable $\mathcal{S}\subseteq\mathcal{T}$ and for all neighboring datasets $\mathbb{D},\mathbb{D}^{\prime}\in\mathcal{D}$.
[…]
Definition 2 ($\ell_{2}$ sensitivity [44]).
The $\ell_{2}$- sensitivity of vector valued function $f(\mathbb{D})$ is $\Delta:=\max_{\mathbb{D},\ \mathbb{D}^{\prime}}\|f(\mathbb{D})-f(\mathbb{D}^{\prime})\|_{2}$, where $\mathbb{D}$ and $\mathbb{D}^{\prime}$ are neighboring datasets.
The $\ell_{2}$ sensitivity of a function gives the upper bound of how much the function can change if one sample at the input is changed. Consequently, it dictates the amount of randomness/perturbation needed at the function’s output to guarantee differential privacy. In other words, it captures the maximum change in the output by changing any one user in the worst-case scenario.
Definition 3 (Gaussian Mechanism [45]).
Let $f:\mathcal{D}\mapsto\mathcal{R}^{D}$ be an arbitrary function with $\ell_{2}$ sensitivity $\Delta$. The Gaussian mechanism with parameter $\tau$ adds noise from $\mathcal{N}(0,\tau^{2})$ to each of the $D$ entries of the output and satisfies $(\epsilon,\delta)$ differential privacy for $\epsilon\in(0,1)$ and $\delta\in(0,1)$, if $\tau\geq\frac{\Delta}{\epsilon}\sqrt{2\log\frac{1.25}{\delta}}$.
Here, $(\epsilon,\delta)$-DP is guaranteed by adding noise drawn form $\mathcal{N}(0,\tau^{2})$ distribution. Note that, there are an infinite number of combinations of $(\epsilon,\delta)$ for a given $\tau^{2}$ [5].
Definition 4 (Rényi Differential Privacy (RDP) [47]).
A randomized algorithm $f$ : $\mathcal{D}\mapsto\mathcal{T}$ is $(\alpha,\epsilon_{r})$-Rényi differentially private if, for any adjacent $\mathbb{D},\ \mathbb{D}^{\prime}\in\mathcal{D}$, the following holds: $D_{\alpha}(\mathcal{A}(\mathbb{D})\ ||\mathcal{A}(\mathbb{D}^{\prime}))\leq\epsilon_{r}$. Here, $D_{\alpha}(P(x)||Q(x))=\frac{1}{\alpha-1}\log\mathbb{E}_{x\sim Q}\Big{(}\frac{P(x)}{Q(x)}\Big{)}^{\alpha}$ and $P(x)$ and $Q(x)$ are probability density functions defined on $\mathcal{T}$.
Proposition 1 (From RDP to DP [47]).
If $f$ is an $(\alpha,\epsilon_{r})$-RDP mechanism, it also satisfies $(\epsilon_{r}+\frac{\log 1/\delta}{\alpha-1},\delta)$-differential privacy for any $10<\delta<$ 1.
Proposition 2 (Composition of RDP [47]).
Let $f_{1}:\mathcal{D}\rightarrow\mathcal{R}_{1}$ be $(\alpha,\epsilon_{1})$-RDP and $f_{2}:{\mathcal{R}_{1}}\times\mathcal{D}\rightarrow\mathcal{R}_{2}$ be $(\alpha,\epsilon_{2})$-RDP, then the mechanism defined as $(X_{1},X_{2})$, where $X_{1}\sim f_{1}(\mathbb{D})$ and $X_{2}\sim f_{2}(X_{1},\mathbb{D})$ satisfies $(\alpha,\epsilon_{1}+\epsilon_{2})$-RDP.
Proposition 3 (RDP and Gaussian Mechanism [47]).
If $f$ has $\ell_{2}$ sensitivity 1, then the Gaussian mechanism $\mathcal{G}_{\sigma}f(\mathbb{D})=f(\mathbb{D})+e$ where $e\sim\mathcal{N}(0,\sigma^{2})$ satisfies $(\alpha,\frac{\alpha}{2\sigma^{2}})$-RDP. Also, a composition of $T$ such Gaussian mechanisms satisfies $(\alpha,\frac{\alpha T}{2\sigma^{2}})$-RDP.